A cylinder of mass M and radius R is rotting down
Answer #1

for the yoyo , force equation in vertical direction is given as

mg - T = ma           

T = mg - ma                                     eq-1

Torque equation is given as

T r = I a/r

(mg - ma) = (mr2) (a/r2)

2 ma = mg

a = g/2

a = 4.9